OpenStudy (anonymous):

Special Products and Factoring.... Please Help... Thankyou :) Pls. See the answer section...

OpenStudy (anonymous):

\[(X ^{n} -4 ) (X ^{n} -5 )\]

OpenStudy (anonymous):

(x^n -4) (x^n - 5)

OpenStudy (anonymous):

x^n=m (m-5)(m-4)=m^2-9m+20 x^2n-9x^n+20

OpenStudy (anonymous):

Why is x^n =m?

OpenStudy (anonymous):

we don't need it.you can find the answer with x^n too.I used it to simplify the question.

OpenStudy (anonymous):

Oh Okay then.. Thankyou Verymuch! :D